a curve has the equation y=3x^2=5x+4

1- find the gradient of the curve at the point (2,6)

2- find the coordinates of the point on the curve where the gradient is -2


1
Expert's answer
2021-05-07T10:20:55-0400

1) "f'(x)=6x-5."

"f'(2)=12-5=7."


2) "f'(x)=-2."

"6x-5=-2."

"x=\frac{1}{2}".

"y=f(\frac{1}{2})=\frac{3}{4}-\frac{5}{2}+4=\frac{9}{4}."

Point on the curve: "(\frac{1}{2},\frac{9}{4})."